OpenStudy (anonymous):

Solve the inequality –2(x – 3) ≥ 5 – (x + 3)

OpenStudy (richyw):

you can treat it like an equal sign, but you have to be careful because it is "directional". So like \(x=9\) is the same as \(9=x\), but \(x<9\) is different that \(9<x\). \[-2(x – 3)\leq 5-(x + 3)\]\[-2(x-3)+(x+3)\leq5\]\[-2x+6+x+3\leq 5\]\[-x+9\leq5\]\[-x\leq-4\]\[x\leq4\]

OpenStudy (anonymous):

ok thankyou